OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation. 3x + 2y = -21 x - 3y = 4 Choose one answer. a. 3(3y + 4) + 2y = -21 b. 3(-3y + 4) + 2y = -21 c. 3x + 2(3y + 4) = -21 d. 3x + 2(-3y + 4) = -21 .

OpenStudy (psymon):

So what we want to do is choose one of the two equations and then isolate one of the variables in that equation. Basically, get x or y by itself on one side of the equal sign. Given your two equations, it looks easiest to do this with the second one. So for the second one, we can get x all by itself by adding 3y to both sides. That will give us this: x = 4 + 3y Now we take this new found value of x and plug it into the other equation, replacing x with this value. Doing that looks like this: 3(4+3y) + 2y = -21. The x was completely replaced with the 4+3y value and put in parenthesis because we want this to be multiplication. So now I need to multiply this out and solve for y. This gives me: 12 + 9y + 2y = -21 12 + 11y = -21 11y = -33 y = -3 So this is how you would go about solving for one of the variables. Now that I know what y is, I can substitute this into either equation and solve for x. The second one looks like the easiest to do this with, so Ill replace y with -3 in the second one. That'll give me: x-3(-3) = 4 x + 9 = 4 x = -5 So that would be our x and y value and the point of intersection for this system, (-5,-3)